# Bridging the Gap: Towards an Expanded Toolkit for ML-Supported Decision-Making in the Public Sector

### v2

@@ 3.1.2 Uncertainty Quantification
In recent years, interest has grown in conformal prediction as a distribution-free and model-agnostic approach to uncertainty quantification for ML models. These characteristics make conformal prediction particularly appealing in many practical scenarios, as no specific assumptions on the model are required, enabling easy implementation for any arbitrary ML model. Instead of providing a point prediction, conformal prediction constructs a set of plausible predictions with respect to a chosen significance level (Vovk et al., , 2022; Angelopoulos and Bates, , 2021; Papadopoulos et al., , 2002). A larger conformal set indicates higher uncertainty in the model’s predictions. However, conformal prediction requires splitting the data into a training set and an additional holdout dataset, known as the calibration set. While this approach enhances model robustness, it comes at the expense of reduced statistical efficiency due to the decreased amount of data available for training. Alternatively, full conformal prediction does not necessitate dividing the data but is usually computationally more demanding (Angelopoulos and Bates, , 2021). Conformal predictions relies on exchangeable data, which can not be guaranteed in scenarios involving distribution shifts. However, efforts have been made to extend conformal prediction for such situations, such as making use of weighting methods akin to those discussed in the context of unsupervised domain adaptation (Tibshirani et al., , 2019; Barber et al., , 2023; Gibbs and Candes, , 2021).
As mentioned, uncertainty estimates play a significant role in facilitating cooperative interaction between human users and models, particularly for high-stakes decision-making prevalent in the public sector. For example, Straitouri and Rodriguez, (2023) propose a decision-support framework that makes use of conformal prediction to improve the cooperation between experts and the ML model. Their modeling framework restricts domain experts to choose their prediction from a set of plausible predictions generated by the model, resulting in better performance than relying on the model or the human expert alone.
@@ 3.1.3 Multi-Objective Optimization
In the context of risk prediction, predictive modeling and decision-making are separated into two distinct steps (Elmachtoub and Grigas, , 2022; Kuppler et al., , 2022). Initially, a prediction is generated that subsequently gets used to inform a downstream allocation problem. In current ADM systems practice, multi-objective optimization is rarely employed. Typically, problems are cast as single-objective constrained optimization tasks, like finding the best allocation within budget constraints. For example, if each individual intervention costs the same, a decision-maker can simply intervene on the top-$k$ ranked individuals.
When more complex constraints, different decision criteria and multiple predictions come into play, a conventional approach for formalizing the decision step involves the creation of a scalar utility function that linearly combines different objectives into a weighted sum (Keeney and Raiffa, , 1993). For instance, stakeholders might construct a unified risk score out of multiple criteria that is subsequently employed to prioritize the allocation of resources. However, constructing a joint utility function can be tricky (Das and Dennis, , 1997), as stakeholders often struggle to determine how to exactly weigh different objectives (Hayes et al., , 2022; Roijers et al., , 2013). This difficulty is especially pronounced in risk prediction, where the link between predictions and the expected utility is often not entirely specified. A predicted risk score may only allow for a prioritization of individuals while the exact size of the individual utilities remains unknown. For example, in scenarios where intervention costs vary significantly by individual it might be important to compare the exact magnitude of the guiding utility for each individual intervention, making it problematic if only a ranked list is available. Boutilier, (2013) offers further insights into techniques for eliciting utility functions and preferences from decision-makers.
@@ 3.2 Counterfactual Prediction
The primary goal of any ADM system is to guide decision-making by recommending a particular course of action. Making such recommendations effectively often involves counterfactual modeling. While non-causal risk predictions are often used as proxies for relevant counterfactual outcomes, they risk being significantly biased, making a more principled approach involving explicit causal modeling preferable. However, as outlined in Section 2.3, a common threat to the validity of causal models is confounding, requiring external assumptions and historical data on intervention assignment to address. This challenge requires careful case-by-case analysis by the model developer and may limit the possibility to make use of counterfactual estimates for decision-making in certain application contexts.
@@
The potential outcomes framework (Rubin, , 1974) is a prominent approach for framing causal questions. In a binary intervention scenario $\mathcal{T}=\{0,1\}$, it denotes two potential outcomes $(Y_{i}(0),Y_{i}(1))$ for an individual $i$. These outcomes represent the two possible observable outcomes: no intervention ($T_{i}=0$) and an intervention ($T_{i}=1$). The individual treatment effect $\tau_{i}$ is then defined as the difference between the potential outcomes $\tau_{i}=Y_{i}(1)-Y_{i}(0)$. Estimating potential outcomes and treatment effects from observational data $\mathcal{D}=\{(X_{i},T_{i},Y_{i})\}_{i=1}^{n}$ is challenging, as it is usually only possible to observe one outcome $Y_{i}=(1-T_{i})Y_{i}(0)+T_{i}Y_{i}(1)$ for each individual (Künzel et al., , 2019). Consequently, it is common to estimate the expected potential outcomes $\mu_{t}(x)=\mathbb{E}[Y(t)|X=x]$ and conditional average treatment effect (CATE) $\tau(x)=\mathbb{E}[Y(1)-Y(0)|X=x]$ for a given covariate vector $X=x$ (Vegetabile, , 2021; Künzel et al., , 2019).
@@
To link the CATE with a statistical estimand, a set of untestable assumptions is required (Künzel et al., , 2019; Caron et al., , 2022; Johansson et al., , 2022). Unconfoundedness $(Y(0),Y(1))\perp\!\!\!\perp T|X$, requires that potential outcomes are conditionally independent of treatment assignment. Positivity guarantees nonzero propensity scores $10<P(T=1|X=x)<$ 1 for all confounders $x\in\mathcal{X}$, meaning that treatment assignment is not fully deterministic. Finally, Stable Unit Treatment Value Assumption (SUTVA) assumes that the outcome of one individual is not affected by the interventions others received, and that there are no different versions of a specific treatment. Under these assumptions it becomes in principle possible to infer $\mu_{t}(x)$ and $\tau(x)$ from observational data (Caron et al., , 2022). Ensuring these assumptions can be difficult, with unmeasured confounding posing a significant risk when aiming for valid counterfactual predictions. We refer to Appendix A for an overview of relevant ML-based CATE estimation methods.
@@
While the individual treatment effect seems like a natural choice for determining the optimal allocation, there exist various scenarios in which estimating only one expected potential outcome $\mu_{t}(x)$ may be sufficient to inform the decision-making process Dickerman and Hernán, (2020). These outcomes could, for example, represent the likelihood of abuse if a hotline call is not followed up (Coston et al., , 2020) or the risk of death of a patient if no heart transplant is performed (Van Geloven et al., , 2020; Dickerman and Hernán, , 2020). Models that provide such risk assessments align well with allocation principles informed by need-based criteria by identifying individuals at high risk of adverse outcomes. For example, when screening phone calls for potential child maltreatment (Vaithianathan et al., , 2019), there exists a moral and legal obligation to investigate high risk cases, regardless of the investigation’s likelihood of success (Coston et al., , 2020; Chouldechova et al., , 2018). Additionally, in scenarios where one potential outcome is trivially known, only one outcome needs to be estimated. For instance, in judicial bail prediction, individuals for whom bail was denied cannot re-offend before trial (Lakkaraju et al., , 2017).
@@ 3.2.1 Distribution Shifts, Selection Bias and Label Bias
The challenge of handling distribution shifts is strongly related to causal estimation, as illustrated by Johansson et al., (2016). For example, predicting counterfactual outcomes under no unmeasured confounding corresponds to unsupervised domain adaptation under covariate shift (Johansson et al., , 2022). This is because past decision-making policies often lead to a difference in covariate distribution between the treatment groups and the distribution of the overall population. Several approaches for dealing with shifts when performing CATE estimation have been proposed (Johansson et al., , 2016). For example, Kuzmanovic et al., (2023) study the problem of inferring CATE in settings in which treatment information is missing for some individuals, a challenge they frame as a covariate shift problem.
In many practical settings, approaches geared towards guaranteeing robustness to unknown distribution shifts (Jeong and Namkoong, , 2020) may be particularly relevant, as it can be difficult to anticipate the target population and relevant subpopulations in all possible deployment environments. To tackle this challenge, Zhou et al., (2023) introduce an approach for learning robust CATE estimates under unknown external covariate shifts. They achieve this by employing a boosting-style post-processing routine to generate a multi-accurate predictor (Kim et al., , 2019), enabling unbiased predictions in a new deployment setting.
@@ 3.2.3 Multi-Objective Optimization
Efforts have been made to explicitly integrate CATE estimation and prescriptive optimization within an unified framework, typically with a focus on budget-constrained optimization problems (Tu et al., , 2021; Ai et al., , 2022). Formulating such optimization problems is generally made easier when the expected net benefit can be easily defined, such as maximizing net revenue when allocating tax audits within a fixed budget (Black et al., , 2022). Similarly, McFowland III et al., (2021) present a prescriptive analytics framework that combines randomized experiments, CATE estimation and a subsequent constrained optimization problem to identify the profit-maximizing allocation policy. Crucially, the expected cost of an intervention may not necessarily be known, potentially requiring a separate estimation process. Unlike in the case of generic prediction, only a few studies have attempted to integrate constraints directly into the counterfactual estimation process. Notable examples include (Kim and Zubizarreta, , 2023) for CATE estimation and (Mishler et al., , 2021) for counterfactual risk prediction under fairness constraints.
@@ 3.3 Policy Learning
The ADM approaches discussed so far entail a two-step process: initially estimating individual outcomes, such as the CATE, and subsequently using these estimates to determine an optimal downstream allocation considering external constraints. However, this means that the target of estimation is only indirectly linked to the underlying policy objective, as an improved prediction may not necessarily enhance the utility of the resulting allocation policy (Perdomo, , 2023). While perfect predictions could in theory lead to optimal decision-making, in practice it may be easier and more practical to estimate the allocation policy directly (Elmachtoub and Grigas, , 2022; Fernández-Loría and Provost, 2022a, ). A wide range of methods have been proposed to optimize the aggregated utility, emphasizing that the primary goal of deploying a statistical targeting system is not accurate prediction alone.
@@
More specifically, the target of estimation becomes the allocation policy $\pi:\mathcal{X}\rightarrow[0,1]$, directly mapping individual covariates $X_{i}$ to the likelihood of intervening. Learning optimal assignment rules has been studied across different disciplines, such as statistical decision theory, economics and operations research (Manski, , 2004; Kitagawa and Tetenov, , 2018; Elmachtoub and Grigas, , 2022). This paper specifically highlights recent literature focusing on learning optimal policies from past observational data using ML methods (Athey and Wager, , 2021; Kallus, , 2018; Hatamyar and Kreif, , 2023; Luedtke and van der Laan, , 2016). Here, we are concerned with off-policy learning, given that on-policy learning may not be suitable for high-stakes settings in the public sector where active experimentation with different decision policies is not possible.
@@
Off-Policy learning usually involves optimizing over a class of policies $\pi\in\Pi$ by defining the aggregated utility of a proposed policy $V(\pi)=\mathbb{E}[\pi(x)Y(0)+(1-\pi(x))Y(1)]$ as the overall expected outcome if the policy were deployed (Athey and Wager, , 2021). Finding the optimal policy corresponds to identifying the policy that maximizes the utility, i.e. $\hat{\pi}=\underset{\pi\in\Pi}{\operatorname*{arg\,max}}\ \hat{V}(\pi)$. When estimating the policy value from observational data, we rely on the same assumptions as those used in CATE estimation to ensure identification, such as unconfoundedness. We refer to Appendix B for an overview of off-policy learning methods.
@@
Adopting a population-level perspective and directly optimizing for the best allocation policy can come with several benefits. Firstly, such an approach may aid in the natural integration of downstream constraints, for example by limiting the class of policies $\Pi$ under consideration to these that can feasibly be implemented. For instance, policy learning enables the exclusion of decision policies that make use of specific covariates (Kallus, , 2021; Athey and Wager, , 2021), such as sensitive attributes like race and gender, or features susceptible to individual manipulation potentially leading to distribution shifts after deployment. While these variables may be required to address confounding during estimation, we can ensure that the decision-making does not rely on them by constraining the class of allowed policies. Secondly, estimating and optimizing the policy value $V(\pi)$ for a constrained set of policies is a distinct and potentially easier estimation task compared to predicting individual-level treatment effects (Kallus, , 2021; Lechner, , 2023). In general, precise estimation of individual-level outcomes may not always be a prerequisite for determining the optimal policy, as an erroneous prediction may not necessarily lead to an erroneous decision (Fernández-Loría and Provost, 2022a, ).
@@
The feasibility of employing policy learning also depends on whether the goal of the modeling process is a fully automated decision system or providing recommendations to human decision-makers. As discussed in Section 2.5, fully formalizing the connection between model output and decision-making can be challenging due to the involvement of human decision-makers who may want to integrate external information and can overrule the model’s recommendation (Shalit, , 2022). This complication adds nuance to the discussion about choosing the most appropriate target estimand and modeling framework. For example, a human decision-maker might find a CATE estimate more trustworthy and more suitable for individual decision-making, as opposed to a fully defined allocation policy (Coston et al., , 2020). Conversely, a well-defined policy class could also be restricted to policies that can be easily interpreted by users, such as decision trees, as seen in (Athey and Wager, , 2021).

### v3

@@ 3.1.2 Uncertainty Quantification
In recent years, interest has grown in conformal prediction as a distribution-free and model-agnostic approach to uncertainty quantification for ML models. These characteristics make conformal prediction particularly appealing in many practical scenarios, as no specific assumptions on the model are required, enabling easy implementation for any arbitrary ML model. Instead of providing a point prediction, conformal prediction constructs a set of plausible predictions with respect to a chosen significance level (Vovk et al., , 2022; Angelopoulos and Bates, , 2021; Papadopoulos et al., 2002a, ). A larger conformal set indicates higher uncertainty in the model’s predictions. Conformal prediction requires splitting the data into a training set and an additional holdout dataset, known as the calibration set. Alternatively, full conformal prediction does not necessitate dividing the data but is usually computationally more demanding (Angelopoulos and Bates, , 2021). Conformal prediction relies on exchangeable data, which can not be guaranteed in scenarios involving distribution shifts. However, efforts have been made to extend conformal prediction for such situations, such as making use of weighting methods akin to those discussed in the context of unsupervised domain adaptation (Tibshirani et al., , 2019; Barber et al., , 2023; Gibbs and Candes, , 2021).
As mentioned, uncertainty estimates play a significant role in facilitating cooperative interaction between human users and models, particularly for high-stakes decision-making prevalent in the public sector. For example, Straitouri and Gomez Rodriguez, (2024) propose a decision-support framework that makes use of conformal prediction to improve the cooperation between experts and the ML model. Their modeling framework restricts domain experts to choose their prediction from a set of plausible predictions generated by the model, resulting in better performance than relying on the model or the human expert alone.
@@ 3.1.3 Multi-Objective Optimization
In the context of risk prediction, predictive modeling and decision-making are separated into two distinct steps (Elmachtoub and Grigas, , 2022; Kuppler et al., , 2022). Initially, a prediction is generated that subsequently gets used to inform a downstream allocation problem. In current ADM systems practice, multi-objective optimization is rarely employed. Typically, problems are cast as single-objective constrained optimization tasks, like finding the best allocation within budget constraints. When more complex constraints, different decision criteria and multiple predictions come into play, a conventional approach for formalizing the decision step involves the creation of a scalar utility function that linearly combines different objectives into a weighted sum (Keeney and Raiffa, , 1993). For instance, stakeholders might construct a unified risk score out of multiple criteria that is subsequently employed to prioritize the allocation of resources. However, constructing a joint utility function can be tricky (Das and Dennis, , 1997), as stakeholders often struggle to determine how to exactly weigh different objectives (Hayes et al., , 2022; Roijers et al., , 2013; Boutilier, , 2013). This difficulty is especially pronounced in risk prediction, where the link between predictions and the expected utility is often not entirely specified. A predicted risk score may only allow for a prioritization of individuals while the exact size of the individual utilities remains unknown. For example, in scenarios where intervention costs vary significantly by individual it might be important to compare the exact magnitude of the guiding utility for each individual intervention, making it problematic if only a ranked list is available.
@@ 3.2 Counterfactual Prediction
The primary goal of any ADM system is to guide decision-making by recommending a particular course of action. Making such recommendations effectively will often involve counterfactual modeling. While non-causal risk predictions can be used as proxies for relevant counterfactual outcomes, they risk being significantly biased, potentially making a more principled approach involving explicit causal modeling preferable. However, as outlined in Section 2.3, a common threat to the validity of causal models is confounding, requiring external assumptions and historical data on intervention assignment to address. This challenge requires careful case-by-case analysis by the model developer and may limit the possibility to make use of counterfactual estimates for decision-making in certain application contexts.
@@
The potential outcomes framework (Rubin, , 1974) is a prominent approach for framing causal questions. In a binary intervention scenario $\mathcal{T}=\{0,1\}$, it denotes two potential outcomes $(Y_{i}(0),Y_{i}(1))$ for an individual $i$. These outcomes represent the two possible observable outcomes: no intervention ($T_{i}=0$) and an intervention ($T_{i}=1$). The individual treatment effect $\tau_{i}$ is then defined as the difference between the potential outcomes $\tau_{i}=Y_{i}(1)-Y_{i}(0)$. Estimating potential outcomes and treatment effects from observational data $\mathcal{D}=\{(X_{i},T_{i},Y_{i})\}_{i=1}^{n}$ is challenging, as it is usually only possible to observe one outcome $Y_{i}=(1-T_{i})Y_{i}(0)+T_{i}Y_{i}(1)$ for each individual (Künzel et al., , 2019). Consequently, it is common to estimate the expected potential outcomes $\mu_{t}(x)=\mathbb{E}[Y(t)|X=x]$ and conditional average treatment effect (CATE) $\tau(x)=\mathbb{E}[Y(1)-Y(0)|X=x]$ for a given covariate vector $X=x$ (Vegetabile, , 2021; Künzel et al., , 2019). We refer to Appendix B for an overview of relevant ML-based CATE estimation methods.
@@
To link the CATE with a statistical estimand, a set of untestable assumptions is required (Künzel et al., , 2019; Caron et al., , 2022; Johansson et al., , 2022). Unconfoundedness $(Y(0),Y(1))\perp\!\!\!\perp T|X$, requires that potential outcomes are conditionally independent of treatment assignment. Positivity guarantees nonzero propensity scores $10<P(T=1|X=x)<$ 1 for all confounders $x\in\mathcal{X}$, meaning that treatment assignment is not fully deterministic. Finally, Stable Unit Treatment Value Assumption (SUTVA) assumes that the outcome of one individual is not affected by the interventions others received, and that there are no different versions of a specific treatment. Under these assumptions it becomes in principle possible to infer $\mu_{t}(x)$ and $\tau(x)$ from observational data (Caron et al., , 2022). Ensuring these assumptions can be difficult, with unmeasured confounding posing a significant risk when aiming for valid counterfactual predictions.
@@
While the individual treatment effect seems like a natural choice for determining the optimal allocation, there exist various scenarios in which estimating only one expected potential outcome $\mu_{t}(x)$ may be sufficient to inform the decision-making process (Dickerman and Hernán, , 2020). These outcomes could, for example, represent the likelihood of abuse if a hotline call is not followed up (Coston et al., , 2020) or the risk of death of a patient if no heart transplant is performed (Van Geloven et al., , 2020; Dickerman and Hernán, , 2020). Models that provide such risk assessments align well with allocation principles informed by need-based criteria by identifying individuals at high risk of adverse outcomes. For example, when screening phone calls for potential child maltreatment (Vaithianathan et al., , 2019), there exists a moral and legal obligation to investigate high risk cases, regardless of the investigation’s likelihood of success (Coston et al., , 2020; Chouldechova et al., , 2018). Additionally, in scenarios where one potential outcome is trivially known, only one outcome needs to be estimated. For instance, in judicial bail prediction, individuals for whom bail was denied cannot re-offend before trial (Lakkaraju et al., , 2017).
@@ 3.2.1 Distribution Shifts, Selection Bias and Label Bias
The challenge of handling distribution shifts is strongly related to causal estimation, as illustrated by Johansson et al., (2016). For example, predicting counterfactual outcomes under no unmeasured confounding corresponds to unsupervised domain adaptation under covariate shift (Johansson et al., , 2022). This is because past decision-making policies often lead to a difference in covariate distribution between the treatment groups and the distribution of the overall population. Several approaches for dealing with shifts when performing CATE estimation have been proposed (Johansson et al., , 2016; Shalit et al., , 2017; Assaad et al., , 2021). For example, Kuzmanovic et al., (2023) study the problem of inferring CATE in settings in which treatment information is missing for some individuals, a challenge they frame as a covariate shift problem.
In many practical settings, approaches geared towards guaranteeing robustness to unknown distribution shifts (Jeong and Namkoong, , 2020) may be particularly relevant, as it can be difficult to anticipate the target population and relevant subpopulations in all possible deployment environments. To tackle this challenge, Kern et al., (2024) introduce an approach for learning robust CATE estimates under unknown external covariate shifts. They achieve this by employing a boosting-style post-processing routine to generate a multi-accurate predictor (Kim et al., , 2019), enabling unbiased predictions in a new deployment setting.
@@ 3.2.3 Multi-Objective Optimization
Efforts have been made to explicitly integrate CATE estimation and prescriptive optimization within an unified framework, typically with a focus on budget-constrained optimization problems (Tu et al., , 2021; Ai et al., , 2022). Formulating such optimization problems is generally made easier when the expected net benefit can be easily defined, such as maximizing net revenue when allocating tax audits within a fixed budget (Black et al., , 2022). Similarly, McFowland III et al., (2021) present a prescriptive analytics framework that combines randomized experiments, CATE estimation and a subsequent constrained optimization problem to identify the profit-maximizing allocation policy. Crucially, the expected cost of an intervention may not necessarily be known, potentially requiring a separate estimation process. Unlike in the case of generic prediction, only a few studies have attempted to integrate constraints directly into the counterfactual estimation process. Notable examples include Kim and Zubizarreta, (2023) for CATE estimation and Mishler et al., (2021) for counterfactual risk prediction under fairness constraints.
@@ 3.3 Policy Learning
The ADM approaches discussed so far entail a two-step process: initially estimating individual outcomes, such as the CATE, and subsequently using these estimates to determine an optimal downstream allocation considering external constraints. However, this means that the target of estimation is only indirectly linked to the underlying policy objective, as an improved prediction may not necessarily enhance the utility of the resulting allocation policy (Perdomo, , 2024). While perfect predictions could in theory lead to optimal decision-making, in practice it may sometimes more practical to estimate the allocation policy directly (Elmachtoub and Grigas, , 2022; Fernández-Loría and Provost, 2022a, ). A wide range of methods have been proposed to optimize the aggregated utility, emphasizing that the primary goal of deploying a statistical targeting system is not accurate prediction alone.
@@
More specifically, the target of estimation becomes the allocation policy $\pi:\mathcal{X}\rightarrow\{0,1\}$, directly mapping individual covariates $X_{i}$ to an intervention. Learning optimal assignment rules has been studied across different disciplines, such as statistical decision theory, economics and operations research (Manski, , 2004; Kitagawa and Tetenov, , 2018; Elmachtoub and Grigas, , 2022). This paper specifically highlights recent literature focusing on learning optimal policies from past observational data using ML methods (Athey and Wager, , 2021; Kallus, , 2018; Hatamyar and Kreif, , 2023; Luedtke and van der Laan, , 2016). Here, we are concerned with off-policy learning, given that on-policy learning may not be suitable for high-stakes settings in the public sector where active experimentation with different decision policies is not possible.
@@
Off-Policy learning usually involves optimizing over a class of policies $\pi\in\Pi$ by defining the aggregated utility of a proposed policy $V(\pi)=\mathbb{E}[Y(\pi(x))]$ as the overall expected outcome if the policy were deployed (Athey and Wager, , 2021). Finding the optimal policy corresponds to identifying the policy that maximizes the utility, i.e. $\hat{\pi}=\underset{\pi\in\Pi}{\operatorname*{arg\,max}}\ \hat{V}(\pi)$. When estimating the policy value from observational data, we rely on the same assumptions as those used in CATE estimation to ensure identification, such as unconfoundedness. We refer to Appendix C for an overview of off-policy learning methods.
@@
Adopting a population-level perspective and directly optimizing for the best allocation policy can come with several benefits. First, such an approach may aid in the natural integration of downstream constraints, for example by limiting the class of policies $\Pi$ under consideration to these that can feasibly be implemented. For instance, policy learning enables the exclusion of decision policies that make use of specific covariates (Kallus, , 2021; Athey and Wager, , 2021), such as sensitive attributes like race and gender, or features susceptible to individual manipulation potentially leading to distribution shifts after deployment. While these variables may be required to address confounding during estimation, we can ensure that the decision-making does not rely on them by constraining the class of allowed policies. Second, estimating and optimizing the policy value $V(\pi)$ for a constrained set of policies is a distinct and potentially easier estimation task compared to predicting individual-level treatment effects (Kallus, , 2021; Lechner, , 2023). In general, precise estimation of individual-level outcomes may not always be a prerequisite for determining the optimal policy, as an erroneous prediction may not necessarily lead to an erroneous decision (Fernández-Loría and Provost, 2022a, ).
@@
The feasibility of employing policy learning also depends on whether the goal of the modeling process is a fully automated decision system or providing recommendations to human decision-makers. As discussed in Section 2.5, fully formalizing the connection between model output and decision-making can be challenging due to the involvement of human decision-makers who may want to integrate external information and can overrule the model’s recommendation (Shalit, , 2022). This complication adds nuance to the discussion about choosing the most appropriate target estimand and modeling framework. For example, a human decision-maker might find a CATE estimate more trustworthy and more suitable for individual decision-making, as opposed to a fully defined allocation policy (Coston et al., , 2020). Conversely, a well-defined policy class could also be restricted to policies that can be easily interpreted by users, such as decision trees (Athey and Wager, , 2021).